OpenStudy (lgbasallote):

quick question... \[\frac{x^2}{x^2 - 4}- \frac{x+1}{x+2}\] LCD is \(x^2 - 4\) or \(x+2\)??

OpenStudy (lgbasallote):

thanks :D

OpenStudy (auctoratrox):

the LCD is x^2 - 4 because x+2 and x^2 - 4 divide it cleanly.

OpenStudy (lgbasallote):

im confused o.O how do you know the LCD?

OpenStudy (auctoratrox):

the LCD is what we learned when we were taught how to add/subtract fractions with different denominators.

OpenStudy (shubhamsrg):

x^2 - 4 --> isnt it the LCM ?

OpenStudy (auctoratrox):

in this case, the LCM and the LCD are the same

OpenStudy (lgbasallote):

uhmm but how do you find the LCD of rational expressions?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow {a \over b} \pm {c \over d} = {ad \pm cb \over bd} }\)

OpenStudy (auctoratrox):

we must have the same denominator if we want to add/subtract numerators, so multiply the fraction on the right by (x-2)/(x-2) to get the LCD

OpenStudy (lgbasallote):

so shouldnt it be (x^2-4)(x+2)??

OpenStudy (lgbasallote):

@ParthKohli

OpenStudy (auctoratrox):

no, because x^2 - 4 is a difference of squares, = (x-2)(x+2)

OpenStudy (lgbasallote):

why x-2?

OpenStudy (lgbasallote):

i mean when you multiply the right by x-2/x-2

OpenStudy (auctoratrox):

we're essentially multiplying the fraction on the right by 1 to get the same denominator as the fraction on the left

OpenStudy (lgbasallote):

but why x-2?

OpenStudy (auctoratrox):

because (x+2)(x-2) = the denominator of the fraction on the left

OpenStudy (lgbasallote):

ohhh we multiply something to make it look like that so x-2

OpenStudy (lgbasallote):

ok thanks ^_^

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x^2 - 4 = x^2 - 2^2 = (x + 2)(x - 2) }\)